2. Suppose that a fluid has density p (function of space and time) and velocity v with no sources or sinks. (A) Show that the rate of change of the mass m of the fluid contained in a region Q is dm d! du. -SITE (B) Suppose further that, if the fluid crosses the boundary, show that dt -- S6w) - neds. (©) Using your results in Q2(A) and 02(B), show that де V. (e) + at = 0. (D) Why the contimity equation for water is given by V. (pv) = 0.
1
Expert's answer
2020-06-11T16:38:44-0400

a) First we consider the mass of a fluid which leaves region "Q" by flowing across an element in "\delta S" of "S" in time "\delta t" . This quantity is exactly that which is contained in a small cylinder of cross section "\delta S" of length "(v*ň).\delta t" hence the rate at which fluid leaves "q" by flowing across element "S\delta" is "\rho(v*ň)\delta S" .

Summing over all such elements"\delta S" , we obtain the rate of flow of fluid coming out of "q" across the entire surface "S" . Hence the rate at which mass flow out of region "q" is "\int \rho(v*ň)\delta S=\int(\rho v).ň\delta S" .

b) By Gauss divergence theorem "\int div(\rho v)\ delta V" , the mass "M" of the fluid possesed by the volume "V" of the fluid is "M=\int\rho\delta V" where "\rho = \rho(x,y,z,t)" with "(x,y,z)" the Cartesian co coordinates of a general point of "q" , a fixed region of space. Since the space coordinates are independent of time"t" ,therefore the rate of increase of mass within "V" is "\delta M\over \delta t""=" "\delta \over\delta t" "=" "(" "\int \rho dV"")" "=" "\int" "\delta \rho \over \delta t""*dV" .

c) The total rate of change of mass is , and thus from (2a) and (2b) we get,

"\int" "\delta \rho \over \delta t" "+diq(\rho v)=0" .

d) When the motion of fluid is steady, then "\delta \rho \over \delta t""=0" . Thus the equation of continuity becomes"div(\rho v)=0" . Which is the same for homogenous and incompressible fluid.